please show steps clearly

Note: solve 1 In the circuit shown below, find ift) the problem in time domain 2500 m2 usov 62512} & H
Answer #1

II) I (0) = 0 A Y (0) = ol. = ou Top t > 0 64 Switch Llesed 25000 vlt) - 2 UF και ζρν 22H 6250 - 1 Node voltage uit = t.di =Ulf) uso tuult) at 25ooi + 2x10x2500x2.di 5 VE) - 450 + 2500 i to.old -450 + 25001 t.o.ol.de 0.oldi +25001 - 450=0 tio.di. de(5 +1000D + 250000) i = 45000 L It consists of General saution and particular antigral. ildiz meurenae + Ipanticular General500 t Crane = (6,+ht) e soo L General solution of iLE) particular intigrala - inz.(t) Co +10000 4250000) i = 45000 (0+500) iis (1-12.20 +3.10 ) ), (0:16) SOO i = 0.18 = 0.18 - particular intigrad p.z noi total current expression . iles a rentie ile* iios con 0 (+0) 412 70-18 * V l) = tilidt = :$10-18 +1+1zt jeres o sout v lt 1 = 4 [0.1 * Se escotiga t. Esok - Sv. escote= 0.18 a ) C, = 6.18 x500 C, = १० V Now 56 1161 - 6.18 + (-6. 18 +906)

answered by: ANURANJAN SARSAM